Workers: Miyo, Simon, Akutsu
- Miyo and Yamamoto reported issues on the digital system.
- It tends to suddenly turn off. Probably the memories make this problem?
- It turned out that the thermometers signals are not passed to the KAGRA's digital system so far. The channels should be registered!
- Continued the test of the heaters and the thermometers as reported in the above URL.
- Made a rack for cloth hangers outside the clean booth.
[Ongoing] Need cares of water springs from anchor bolts for oplev pylons.
[Ongoing] Need to confirm what are the scratch lines made by the measurement by the KEK team last week.
- Complete the calibration of the oplev; the arm length was reported by Simon, so....
- Check the balance of the suspended WAB with some wires and sensors are attached.
- Check how the damping would be now after the replacement of the damping magnets.
- Check the reproducibility of the alignment of the suspension among several modes.
- Prepare calibration of photosensors.
- Check heights of several points.
- Wash the coil in ethanol at Univ. of Toyama...?
- Take care of water spring??
- A tweezer.
- Two burndy connectors, vacuum compatible.
- Chairs in the clean booth
- Cutters and scissors!!
Stages outside the chamber to ease wiping our legs.
Good rail(?) to move the WAB suspension in the cryostat in the main beam direction.
Chain block? to support the load in the chamber.
- A clock to know the time easily.
- M6 screw set.